MARC (marc_d):

Given s=t^2-5t+6 calculate the distance travelled,in m,in the fifth second. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

What are the roots (x intercepts) of y = x^2 - 5x + 6 ?

MARC (marc_d):

x=3 and x=2

jimthompson5910 (jim_thompson5910):

actually, s is often a position function. Is that what your book is saying?

MARC (marc_d):

yess ^

MARC (marc_d):

so,what should we do next?

jimthompson5910 (jim_thompson5910):

one moment

OpenStudy (sshayer):

calculate s5-s4 i.e.,t=5 and t=4

MARC (marc_d):

okay,i think i got it.. ^ Thank you @jim_thompson5910 @sshayer :)

jimthompson5910 (jim_thompson5910):

if you graphed y = x^2-5x+6, you'll find that the min is (2.5, -0.25) the largest y gets on the interval [0,5] is y = 6 the smallest y gets is y = -0.25 the distance between those two y values is what?

MARC (marc_d):

is it 6.25?

jimthompson5910 (jim_thompson5910):

yes so on the first half of the parabola, you go 6.25 and the same for the second half 2*6.25 = 12.5 is the total distance traveled

MARC (marc_d):

okay..